OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . .

OpenStudy (jango_in_dtown):

this is an AP series with 1st term=a=1 and common difference=d=-5 also here n=12 so S_n=(n/2)(2*a+(n-1)*d)=(12/2)(2*1+(12-1)*(-5))=-318

OpenStudy (jango_in_dtown):

@lxoser check the solution and do tell me if you have any problem

OpenStudy (ghostgate):

Well the difference is -5, so your 12 digits would be: 1, -4, -9, -14, -19, -24, -29, -34, -39, -44, -49, -54 Add them together and you get: -318. However @jango_IN_DTOWN is right.

OpenStudy (jango_in_dtown):

@Ghostgate, you method is applicable only when the number of terms is less, or else it will take more time.:)

OpenStudy (ghostgate):

@jango_IN_DTOWN That is true, I remember learning this, however I forgot the equation.